What are the foci of the ellipse with equation (x^2/100)+(y^2/64)=1
a.(0,+or- sqroot of 164)
b.(+or- sqroot of 164,0)
c.(0,+or- 6)
b.(+or- 6,0)
i got b in part guess mode ^_^

Answer by Nate(3500)

(x^2/100)+(y^2/64)=1
The ellipse has a horizontal major axis with a center at (0,0).
a=10
b=8
c^2=a^2-b^2
c^2=100-64
c=6
Foci: (6,0),(-6,0)
Well, you are right. It is the second 'b'!